d) A random sample of size 36 was taken from a population distributed as Nμ,3.92.The value of the sample x was 15.6. i. Find a 90% confidence interval for μ.                                 (5mks)

It is believed that value of μ is 17. Use your confidence interval to comment on this belief.(2mks)

  1. Using the provided information below,

x=33.2 , x2=131.67,  y=30.78 , y2=116.52xy=119.8 , and n=10

  1. Find the sample correlation coefficient r                                                                 (4mks)
  2. Find the co-efficient of determination                                                                    (2mks)
1
Expert's answer
2022-03-21T04:19:30-0400

d) "CI=x±z*\frac{\sigma}{\sqrt{n}}"

The critical value for 90% confidence interval "z=\u00b11.645"

Then "CI=15.6±1.645*\frac{3.92}{\sqrt{36}}"

The lower limit is "15.6-1.07=14.53"

The upper limit is "15.6+1.07=16.67"

The confidence interval is therefore "(14.53,16.67)"


Since the interval does not contain the value of "\mu=17" , we can conclude that the value of the population mean "\mu≠17" .



1)

"r=\frac{n∑xy-∑x∑y}{\sqrt{[n∑x^2-(∑x)^2][[n∑y^2-(∑y)^2]}}"

"=\frac{10(119.8)-33.2(30.78)}{\sqrt{[10(131.667)-33.2^2][10(116.52)-30.78^2}}"

"=\frac{1198-1024.884}{\sqrt{(214.46)(217.7916)}}"

"=\frac{173.116}{\sqrt{46707.58654}}"

"=\frac{173.116}{216.1194}"

"=0.801"


2) "r^2=0.6416"